Project

General

Profile

News

MICE News

Added by Kyberd, Paul about 6 years ago

Report from the Target Group

During running on Wednesday 13 July, the target in ISIS, which had previously
been operating normally, performed two actuations where it was observed
to hit the lower stop rather than returning to "hold" position. Shortly
afterwards, it overshot the hold position and went too high. This
behaviour has not been observed before, and is thought to be due to a
malfunction of the position-measuring system.

The target frame has now been raised so that the target cannot intercept
the beam. Further investigations are in progress, but it appears likely
that the target has jammed in the upper bearing and cannot be
actuated. If this is the case, the target will be out of action until
it can be replaced in the next ISIS shut-down".

Implications of the target failure: Spokesman

This target has successfully performed 571k pulses and it had been
decided in agreement with ISIS to leave it in the ring until it broke
while better design was being developed. Both of these things have now
happened. We are fortunate enough to have at least one design that works
(and an even better one in preparation).

This means that the summer run had to be canceled, something that caused
disappointment -- we all like to see MICE alive and running!
Nevertheless the EMR test, which was the main aim of the run, could be
sufficiently successful in the few days of running starting 29 June. The
EMR will therefore be sent back to University of Geneva.

The baseline schedule is that we re-install a new target before the
planned running in February-March 2012 (full EMR run). If anyone has a
strongly motivated proposal for data taking before then please let the
MOM and Spokesperson know as soon as possible.

There is no reason to believe that this incident will have an impact on
the schedule that was agreed at the collaboration meeting (EMR run in
February 2012, StepIV run in October 2012 to March 2013).

August Bank Holiday at RAL: Friday 26th – Monday 29th August:

· R22 closes at 4pm on Fri 26th Aug and re-opens at 07:30 on Tuesday 30th August
· R1 Coffee Lounge closes at 4pm on Fri 26th Aug and re-opens at 08:30 on Wednesday 31st August
· Electron cafe closes at 3:30pm on Fri 26th Aug and re-opens at 08:30 on Tuesday 30th August
· Breakfast will be available in Ridgeway House for anyone who is at RAL over the weekend

Reduced service on Tuesday 30 August 2011:
· R22 open for breakfast and lunch only, from 07:30 to 16:00 (there will be no dinner service)
· Electron Cafe open, from 08:30 to 15:30
· R1 Coffee Lounge closed

Both beams are down and any guests in staying in Ridgeway will be advised on alternative venues for dinner.

MICE news 19th June

Added by Kyberd, Paul about 6 years ago

MICE News - 19th June 2011

TOF1 has been re-installed at RAL after refurbishing in Milano Bicocca
in the week 13/6-17/6. Unfortunately not all PMTs have been exchanged, as
some to be refurbished are still in Japan at Hamamatsu, due to delays
from the last Tsunami. The EMR support nicely fitted inside the
Rm3 downstream platform. So both TOF2, KL and the first planes of EMR
are in place for the next data taking.

MICE news May 30 2011

Added by Kyberd, Paul about 6 years ago

Mice Special news from the Spokesmouse

MICE Schedule Review

As I announced in the MICE video conference, in preparation for the MICE project board of 28 June, we held 23-25 May an internal schedule review of MICE.

The agenda with presentations can be found here https://indico.cern.ch/conferenceDisplay.py?confId=136454 Notes and the recommendations by external reviewers will be posted asap.

Please note that this is an internal review and that the information contained in the talks is not "MICE-official"!

The review was very useful, was quite interactive, raised a number of important questions and provided many elements. The information thus gathered will allow to re-baseline the schedule, at least to step IV. Here are some (selected) highlights. <p>

Step I
1. The EMR test run in July is confirmed, it has been proposed to delay the full EMR data
taking to the winter period (February-March 2012)

Step IV
1. There is now a schedule and milestones for the spectrometer solenoids, the manpower has been considerably reinforced, we are only waiting for the final decision on the quench protection system to start the clock ticking.
2. The first AFC magnet has been wound and a schedule exists for commissioning at RAL in the fall of this year (2011).
3. the LH2 absorber number one is on its way to RAL from Japan.
4. the LH2 R&D is progressing well in the MICE hall, next milestone in July 2011 with LHe filling, final results by the end of 2011.
5. the solid absorbers are getting defined and the diffuser is in construction at Oxford.
6. and as can be expected there is some head-scratching to do to dispatch efficiently all
these activities between the MICE hall and the R9 building at RAL.
7. Controls and monitoring are taking shape -- this is becoming a real cornerstone and urgency of the experiment, volunteers are more than welcome to join, please get in touch with Pierrick Hanlet.<br>
8. Software organization is evolving to a more structured and controlled environment --
help from every involved MICE is important.

Taking all into account, this will lead to a new schedule of MICE step IV by mid-june, in time for the MPB on 28 June and final approval by MICE at the next collaboration meeting.

STEP V-VI
1. The Coupling coils are the next priority of the LBNL group, who are presently actively
searching a large cryostat to test the first, already wound, full size coil.
2. There was considerable progress in the design of the layout of RF power in the hall.
3. We should work hard to be in position to have achieved data taking to demonstrate ionization cooling before the next long ISIS shut down (8 months) starting in August 2014.

Again, many thanks to all contributors, and in particular to Andy Nichols who masterminded the operation.

New MICE online Group Coordinator

Linda Coney as new MICE online group (MOG) coordinator

It is a pleasure to inform you all MICE that Linda has agreed to take
on the co-ordination of the online group in replacement of
Jean-Sebastien who is leaving MICE for another life in July. Let us take
this opportunity to thank "JSG" for having opened the trail and led the
way for the last 3 years, and the online group for ensuring the success
of data taking for step I. The Online group will remain a cornerstone of
MICE activities until the last muon has made it though MICE step VI.

Everyone is enthusiastic about Linda and we are convinced that all will
push in the same direction to help the group achieve its missions!

MICE News. May 22

Added by Kyberd, Paul about 6 years ago

TOF2-KL
The changes in the TOF2-KL trolley have been made as designed
and it's ready for the EMR integration. <br>
TOF2 and KL detectors have been reinstalled along the beam line
and all PMTs checked.

MICE News 15 May

Added by Kyberd, Paul about 6 years ago

News from Tesla
The winding of the two coils on the bobbin for the first Absorber and Focus Coil Cryostat has been completed at Tesla Engineering.

Cerenkov
D. Sanders and L. Cremaldi will visit RAL in mid-June to work on CKOV detector and interact with the MICE software groups.

Power Status in the MICE Local Control Room
DAQ machines OFF
PPS rack OFF
Target rack OFF
DS rack OFF
Magnet rack OFF
RAID machines OFF
ONREC01 OFF
OPIPC1 & 3 OFF
MICESTORE OFF

Network switches ON UPS on wall power
IOCPC1 ON wall
ONREC02 ON wall
ECSERV2 ON wall
IOCPC2 ON wall
ECSERV ON wall
CAGATEWAY ON wall
OPIPC2 ON wall

Checked rack room and all consumer units isolated OFF
Checked all hall racks and all consumer units isolated OFF

MICE News - 13 March 2011

Added by Tunnell, Christopher over 6 years ago

MICE News - 13 March 2011

Status following Power Supply Problems

The DAQ system has been tested and appears to be fully operational.
All the DAQ pcs behave normally
All VME, NIM and CAMAC crate work normally.
All the VME boards respond normally to the initialization process and
some data has been taken with a pulser trigger.
The logic and distribution of the DAQ Trigger, the DAQ busy, the Start of Spill
and the End of Spill signals appear normal.

That's it for the good news.
Unfortunately, the High Voltage crate supplying the TOFs and KL has been found
seriously damaged and had to be sent back to the manufacturer for repair.
This prevents us for testing the trigger logic system and the amplification/shaping
and discrimination chain for the TOF and KL signals. However, the fact that no
electronics module (VME or NIM) involved in the DAQ system has been damaged is
a strong indication that the probability of finding a damaged module in the
crates is small.
The TOF and KL shapers will be tested manually on Tuesday the 22nd since they
are critical, home made, items.
A test run is scheduled on Wednesday the 23rd with only GVA1, the CKOV and
the PBM detectors read out.

MICE News - 13 February 2011

Added by Tunnell, Christopher over 6 years ago

MICE News - 13 February 2011

GRID
RAL plan to take CASTOR down next Tuesday for upgrades to the disk
servers.
This will change the way CASTOR handles checksumming, so with a bit of
luck it should then become possible to further streamline the data
movement.

Printing in R76
Instructions from Kevin Dunford on printing with the OKI printer in R76
on to three printers in PPD can be found at
<http://micewww.pp.rl.ac.uk/projects/mice/wiki/Printers>

Wireless networks at RAL
Some useful information on the wireless at RAL thanks to Dave Sankey.

You all know and love our ticketing system for WiFi at RAL.
Since before it was even introduced we have been pushing for a system
better suited to long term visitors and over the past few months
there has finally been some useful progress.

In particular, if you are a regular visitor to RAL you can now have a
permanent account.

Thus today, if you look on the RAL wireless network you see three
different SSIDs:
i) STFC-Guest
ii) STFC-Facilities-Users
iii) STFC-Staff

The first of these, STFC-Guest, is the SSID you use if you are an
infrequent visitor to RAL (someone who has not been registered in our
Accommodation system CDR).

The second of these, STFC-Facilities-Users, is the SSID you use if
you are a frequent visitor to RAL who has been registered in CDR and
who has a site access card - the new mechanism.

(The third, STFC-Staff, is intended for members of STFC, where
instead of having to authenticate you are on a captive network from
where you need to VPN into the RAL internal network using your STFC
credentials.)

Now for the details...
STFC-Guest needs little further comment. Each time you come you need
to request a ticket through your nominal RAL supervisor (generally
Rose Hayes) and you use the ticket details to authenticate and gain
access to our visitors network. You can request a ticket for a
suitable period (for example a one month ticket) and you can request
that your ticket details be emailed to you prior to your arrival, so
that you can have access as soon as you arrive.

The STFC-Facilities-Users SSID on the other hand doesn't use tickets.
Instead you have a 'permanent' account, tied to your registration in
CDR, and you use these credentials to authenticate and gain access to
the same visitors network as for STFC-Guest. The advantage of course
is that once set up your credentials remain valid for the duration of
your registration in our system.

The procedure for enabling this is that first you need to have your
status changed from Visitor to Facility User and be assigned a
Federal ID. You then need to have the account enabled and you need
to set a password.

We have now completed the first of these sets of steps for those of
you who are registered in CDR. The last step involves you passing by
our Computing Group to set a password that meets our requirements and
to sign that you accept our Acceptable Use Policy.

You should have thought of a suitable password and read this Policy
before you come. The file
<http://www.ppd.stfc.ac.uk/public/PPFinance/stfc_password.pdf> gives
advice on what is allowed for federal passwords and our Acceptable
Use Policy is available at <http://www.stfc.ac.uk/aup&gt;.

We also have two web sites accessible from outside which allow you to
manage your password (details in the stfc_password.pdf above which
was originally aimed at STFC staff, so the advice to contact your
local IT Service Desk team does not apply).
The first, <https://testpassword.stfc.ac.uk&gt;testpassword,
allows you to compose and test your password and the second, a QUEST server at
https://access.stfc.ac.uk>, allows you, once your account has been
enabled, to change your password (this latter may complain about what
browser you use, but seems fine if you click the Ignore button).
We strongly recommend that you test your password on the testpassword
site where the reasons for failing a password are given, prior to
coming to set your password and accept our AUP. However there are
dictionary tests that this site does not perform so it is just
possible that a password acceptable on testpassword may subsequently
be rejected.

Thanks to Dave for this complete description.

MICE News - 06 February 2011

Added by Tunnell, Christopher over 6 years ago

MICE News - 06 February 2011

MALCOLM ELLIS

Cooee!
Following several enjoyable months as a "kept man", Malcolm has found
a bank willing to let a Physicist loose on the inside and is now working as
a Quantitative Analyst in the Financial Markets and Treasury Risk group
of Australia's oldest bank.
He is currently learning about counterparty credit risk and is working on
improving the performance and adding functionality to a Monte Carlo
system used by the bank to assess risk.
While whatever spare mental capacity may be available is mostly being
directed at the Black-Scholes equation and a long new list of
three letter acronyms, he can still just about remember what a
VSLB is and is looking forward to reading about successes in MICE
in the coming year(s).

MOM ROTA
The MOM rota for 2011 is now complete
A new ISIS schedule including the machine run-ups and MPs is available
Both have been uploaded on the MICO page

RUN PLAN
A draft run plan can also be found there as 'MOMs, ISIS and MICE
activities' -- this will be discussed further at the next MICE
collaboration meeting.

NETWORK
The network switches in the MLCR both appear to be working properly now
- we almost have micenet back and working again!

GRID
the LFC will have a scheduled DB update on Wed. 9th - disruption
to data access via Grid

TOUR
During the last CM held at RAL Henry Nebresnky organised some tours of the Tier 1
computing facility for MICE. Anybody interested on going on such a tour
during CM29 should contact Henry before Wednesday
If there is enough response he will try to organise another one.

(91-100/112)

Also available in: Atom